# GenNBV: Generalizable Next-Best-View Policy for Active 3D Reconstruction

### v1

@@ Abstract
11footnotetext: Corresponding author.22footnotetext: Project website: gennbv.tech
@@ 1 Introduction
To overcome these two challenges, we propose GenNBV, an end-to-end generalizable NBV policy. It has a novel design of action spaces and state embeddings, which allows free space exploration and ensures a large coverage when applying to unseen objects. We firstly extend former limited action space, like a hemisphere, to 5D free space. As shown in Fig. 1, this new space is composed of a position subspace of approximately $20m$ x $20m$ x $10m$ and an omnidirectional heading subspace. This free space enables our agent drone to scan from any viewpoint. Moreover, with much larger action space, our RL-based policy can be easily generalized to building-scale geometries without concerning the potential scale shift from training to evaluation. On the contrary, many hand-crafted designs [25, 17] limit available viewpoints to a small predefined set, and thus cannot generalize if there is drastic change in scales.
@@
To validate the effectiveness of GenNBV, we construct a benchmark for training and evaluation with Houses3K [27] dataset from the NVIDIA Isaac Gym [22] simulator. We further test its generalization ability on novel buildings and object categories from OmniObject3D [44] and Objaverse [7] datasets. Our evaluation metrics quantify the completeness, accuracy, and efficiency for active 3D reconstruction. To consolidate completeness and efficiency into a single number, we also propose a metric, Area Under the Curve (AUC) of the coverage. Compared with heuristic [13], information gain-based [13, 17, 48], and RL-based baselines [28], our method achieves the best result under different metrics. We also provide visualizations to demonstrate the generalizability of GenNBV. A demo video is attached in the appendix to further illustrate our method.
@@ 2 Related Work
Traditional 3D Reconstruction. Photometry and geometry are two crucial aspects of reconstruction evaluation. Neural implicit representation [26, 23] has shown progress in photometric rendering performance but faces challenges like time-consuming optimization and poor generalization, hindering its application in real-time reconstruction problems such as 3D SLAM [38, 40, 24, 52, 32, 14]. Geometry, in contrast, typically corresponds to 3D representations such as point clouds and 3D mesh and is directly related to issues like collision avoidance. Considering these properties, researchers are prone to focus on geometric reconstruction in practical applications and we also follow this stream.
Active 3D Reconstruction. Active 3D reconstruction is a promising field that has not been thoroughly benchmarked yet. The pipeline of active 3D reconstruction frameworks alternates between inferring optimal viewpoints, capturing new data, and updating the rebuilt 3D model. With the optimal sequence of key viewpoints, a continuous planning path can be generated using the classic planners, such as RRT [16] and FMM [35].
Existing works can be pivotally differentiated based on the paradigm of NBV policies: rule-based or learning-based. Rule-based approaches like uncertainty-driven works [13, 17, 48, 30] typically yield the next best view from hand-crafted rules from the scene representation, which tends to overfit specific scenes. Most learning-based policies [5, 3, 28] use a deep reinforcement learning algorithm like PPO [34] to sequentially predict the optimal viewpoints based on observation. They must obtain feedback from task-related rewards such as coverage ratio [28, 10] and optimize during massive iteration with task environment.
@@ 3.2 Generalizable State Embedding
Specifically, we first build two mid-level representations that can better model the relationship between the 3D object and our agent: a 3D geometric representation $F^{G}$ from depth maps and a semantic representation $F^{S}$ from RGB images. Then we encode these mid-level representations and concatenate them with a pose embedding into state embedding $s_{t}$, guiding the NBV policy for subsequent decision-making. The overall encoder for state embedding is shown in Fig. 2, and details of each representation are discussed below.
@@ 3.3 Reward Function and Optimization
We train the end-to-end policy with reinforcement learning (RL) and hence design a reward function to precisely reflect the task objective for 3D reconstruction. The policy is optimized with proximal policy optimization [34] (PPO) for parallelizing sampling.
Reward Function. With the occupancy probability $F^{G}_{t}$ at time step $t$, we can threshold each voxel with an empirical bound to determine if it is occupied. This discrimination process outputs a binary occupancy grid with $\tilde{N}_{t}$ voxels being occupied, which is used to calculate the coverage ratio:
@@
In free-space exploration, we also need to avoid collision. Previous limited-space agents [17, 48, 10] do not consider collision avoidance since their search space, like hemisphere, is by-design safe. Thus, we add a negative reward for collision and terminate the episode if a collision happens or the current coverage ratio is greater than 95%. We also implement a negative reward when the number of captured keyframes is over an empirical threshold to improve the path efficiency.
Policy Optimization. Once the reward function has been specified, the policy can be learned through any off-the-shelf RL algorithm. In this work, we specifically use PPO as thousands of workers can be parallelized to improve the sample efficiency. Specifically, given our parameterized policy $\pi_{\theta}$, the objective of PPO is to maximize the following objective function:
@@ 4 Experiments
InfoGain
†Uncertainty-Guided [17] 83.13% 89.31% 0.44 69.08% 92.85% 0.42
@@
In this section, we conduct experiments on Houses3K [27], OmniObject3D [44], and Objaverse [7]. For our policy and other learning-based policies, we train them on Houses3K training datasets. After this, we evaluate all policies on the Houses3K test dataset and the OmniObject3D dataset for quantifying the in-distribution and out-of-distribution generalizability. Finally, we show the visualization results on three datasets, demonstrating the effectiveness of the proposed method.
@@ 4.1 Experimental Setup
Dataset. We conduct our experiments on Houses3K [27], OmniObject3D [44] and Objaverse 1.0 [7]. Our model is trained on large-scale 3D objects from Houses3K. To validate the effectiveness and generalizability of our model, we evaluate it on unseen 3D objects from Houses3K. For testing the cross-dataset generalizability, we also evaluate it on batches of various objects from OmniObject3D and Objaverse datasets, which include both house and non-house categories.
@@
Category Dataset NBV Policy AUC normal-$\uparrow$ FCR normal-$\uparrow$ Acc. normal-$\downarrow$
@@
Implementation Details. We conduct all experiments in Isaac Gym simulation engine with one NVIDIA Tesla V100 GPU. Our implementation refers to the codebase of Legged Gym [33] and the PPO implementation in Stable Baseline3 [29]. The ground-truth point clouds on objects’ surfaces are sampled by the Poisson Disk sampling method [47] using Open3D API [51]. The Chamfer Distance between the scanned point cloud and the ground-truth point cloud is calculated using PyTorch3D API [31]. Please refer to the Appendix for further details.
@@ 4.2 Performance Comparison
Non-house Generalization In Table 2, we introduce an additional experiment to evaluate the generalizability of our GenNBV trained on Houses3K to non-house categories. The targets include $74$ animals, $43$ trucks and $33$ dinosaurs from OmniObject3D, and a challenging indoor scene (Room 0) from Replica. The number of views is set to 20 for all experiments.
@@ 4.3 Ablation Study
4.4 Qualitative Results
5 Conclusion
This study presents an end-to-end approach for active 3D scene reconstruction, reducing the need for manual intervention. Specifically, the learning-based policy explores how to reconstruct diverse objects in the training stage and thus can generalize to reconstruct unseen objects in a fully autonomous manner. Our controller maneuvers in free space and the next best view selection based on a hybrid scene representation which conveys scene coverage status and thus reconstruction progress. We show the effectiveness of our approach by testing it on multiple datasets including Houses3K, OmniObject3D, and Objaverse. The quantitative and qualitative generalization results on holdout Houses3K test set and cross-domain OmniObject3D house category show that our method outperforms other baselines in terms of reconstruction completeness, efficiency and accuracy. In addition, the experiment conducted on Objaverse shows that the policy trained in single-building settings can even generalize to a complicated outdoor scene.
Supplementary Material
Appendix A Implementation Details
A.1 Occupancy Grid Mapping Algorithm
We implement this algorithm with PyTorch. In particular, we used ray-casting renderers from PyTorch3D [22] to generate and use depth maps as inputs to our model.
. Occupancy Threshold Mean AUC normal-$\uparrow$ Final Coverage Ratio normal-$\uparrow$ Accuracy normal-$\downarrow$ 0.5 (Ours) 87.39% 95.63% 0.37 1.0 84.41% 93.43% 0.41 1.5 84.84% 95.20% 0.39 2.5 54.77% 58.70% 0.85 Value of Constant $C^{\prime}$ Mean AUC normal-$\uparrow$ Final Coverage Ratio normal-$\uparrow$ Accuracy normal-$\downarrow$ 5 66.96% 71.92% 0.49 10 80.75% 92.16% 0.43 20 (Ours) 87.39% 95.63% 0.37 40 86.97% 94.12% 0.40
A.2 Network and Training
Training.In Isaac Gym [22], we set 256 parallel environments for policy training, where each environment corresponds with one building-level object. Once the coverage ratio reaches a threshold (99% in practice), a collision happens or the episode length reaches the maximum threshold (100 steps), the environment is reset and the building to be reconstructed is replaced. Our NBV policy is optimized through over 32 million iterations and using approximately 24-hour training time on a NVIDIA V100 GPU. All networks are randomly initialized and trained end-to-end.
$s_{t}=\mathrm{Linear}(s^{G}_{t};s^{S}_{t};s^{A}_{t}),$ (7)
where all embeddings are 256 vectors.
Our policy network PPO, implemented by Stable Baseline3 [29], is a 3-layer multi-layer perceptron network (MLP). The output of our policy is used to parameterize a distribution over our 5-dimension action space. In this way, the action can be drawn from the stochastic policy $a\sim\pi(\cdot|o_{t})$.
A.3 Baseline Policies
Appendix B Data Preprocessing
Appendix C Additional Experiments
@@
The effectiveness of 2D semantic representations
Below, we add the experiments on Houses3K test set to illustrate the effectiveness of the proposed semantic representation. Note that Scan-RL’s representation uses 6 frames, while our semantic representation only needs 2 frames.
Num. of RGB NBV Policy AUC normal-$\uparrow$ FCR normal-$\uparrow$ Acc. normal-$\downarrow$

### v2

@@ Abstract
11footnotetext: Corresponding author.
@@ 1 Introduction
To overcome these two challenges, we propose GenNBV, an end-to-end generalizable NBV policy. It has a novel design of action spaces and state embeddings, which allows free-space exploration and ensures a large coverage when applying to unseen objects. We firstly extend former limited action space, like a hemisphere, to 5D free space. As shown in Fig. 1, this new space is composed of a position subspace of approximately $20m$ x $20m$ x $10m$ and an omnidirectional heading subspace. This free space enables our agent drone to scan from any viewpoint. Moreover, with a much larger action space, our RL-based policy can be easily generalized to building-scale geometries without concerning the potential scale shift from training to evaluation. On the contrary, many hand-crafted designs [25, 17] limit available viewpoints to a small predefined set, and thus cannot generalize if there is a drastic change in scales.
@@
To validate the effectiveness of GenNBV, we construct a benchmark for training and evaluation with Houses3K [27] dataset from the NVIDIA Isaac Gym [22] simulator. We further test its generalization ability on novel buildings and object categories from OmniObject3D [44] and Objaverse [7] datasets. Our evaluation metrics quantify the completeness, accuracy, and efficiency for active 3D reconstruction. To consolidate completeness and efficiency into a single number, we also propose a metric, the Area Under the Curve (AUC) of the coverage. Compared with heuristic [13], information gain-based [13, 17, 48], and RL-based baselines [28], our method achieves the best result under different metrics. We also provide visualizations to demonstrate the generalizability of GenNBV. A demo video is attached in the appendix to further illustrate our method.
@@ 2 Related Work
Traditional 3D Reconstruction. Photometry and geometry are two crucial aspects of reconstruction evaluation. Neural implicit representations [26, 23] have shown progress in photometric rendering but faces challenges like time-consuming optimization and poor generalization, hindering its application in real-time reconstruction scenarios like 3D SLAM [38, 24, 52, 32]. Geometry, in contrast, typically corresponds to 3D representations such as point clouds and 3D mesh and is directly related to issues like collision avoidance. Considering these properties, researchers tend to focus on geometric reconstruction in practical applications and we also follow this stream.
Active 3D Reconstruction. Active 3D reconstruction is a promising field that has not been thoroughly benchmarked yet. The pipeline of active 3D reconstruction frameworks alternates between inferring optimal viewpoints, capturing new data, and updating the rebuilt 3D model. With the optimal sequence of key viewpoints, a continuous planning path can be generated using the classic planners, such as Fast Marching Method (FMM) [35].
Existing works can be pivotally differentiated based on the paradigm of NBV policies: rule-based or learning-based. Rule-based approaches like uncertainty-driven works [13, 17, 48, 30] typically yield the next best view from hand-crafted rules from scene representations, which tends to overfit specific scenes. Most learning-based policies [5, 3, 28] use a deep reinforcement learning algorithm like PPO [34] to sequentially predict the optimal viewpoints based on observation. They must obtain feedback from task-related rewards such as coverage ratio [28, 10] and optimize during massive iteration with task environment.
@@ 3.2 Generalizable State Embedding
Specifically, we first build two mid-level representations that can better model the relationship between the objects and our agent: a 3D geometric representation $F^{G}$ from depth maps and a semantic representation $F^{S}$ from RGB images. Then we encode these representations and concatenate them with a pose embedding into state embedding $s_{t}$, guiding the NBV policy for subsequent decision-making. The overall encoder for state embedding is shown in Fig. 2, and details of each representation are discussed below.
@@ 3.3 Reward Function and Optimization
We train the end-to-end policy with reinforcement learning (RL) and hence design reward functions to precisely reflect the task objective for 3D reconstruction. The policy is optimized with proximal policy optimization [34] (PPO) for parallelizing sampling.
Reward Functions. With the occupancy probability $F^{G}_{t}$ at time step $t$, we can threshold each voxel with an empirical bound to determine if it is occupied. This discrimination process outputs a binary occupancy grid with $\tilde{N}_{t}$ voxels being occupied, which is used to calculate the coverage ratio:
@@
In free-space exploration, we also need to avoid collision. Previous limited-space agents [17, 48, 10] do not consider collision avoidance since their search space, like hemisphere, is by-design safe. Thus, we add a negative reward for collision and terminate the episode if a collision happens. We also implement a negative reward when the number of captured keyframes is over an empirical threshold to improve the path efficiency.
Policy Optimization. Once the reward functions has been specified, the policy can be learned through any off-the-shelf RL algorithm. In this work, we use PPO as thousands of workers can be parallelized to improve the sample efficiency. Specifically, given our parameterized policy $\pi_{\theta}$, the objective of PPO is to maximize the following function:
@@ 4 Experiments
InfoGain †Uncertainty-Guided [17] 83.13% 89.31% 0.44 69.08% 92.85% 0.42
@@
In this section, we conduct experiments on Houses3K [27], OmniObject3D [44], and Objaverse [7]. For our policy and other learning-based policies, we train them on Houses3K training set. Then we evaluate all policies on the Houses3K test dataset and the OmniObject3D dataset for quantifying the in-distribution and out-of-distribution generalizability. Finally, we show the visualization results on three datasets, demonstrating the effectiveness of the proposed method.
@@ 4.1 Experimental Setup
Dataset. We conduct our experiments on Houses3K [27], OmniObject3D [44], Objaverse 1.0 [7] and Replica [37]. Our model is trained on large-scale 3D objects from Houses3K. To validate the effectiveness and generalizability of our model, we evaluate it on unseen 3D objects from Houses3K and further on batches of unseen cross-domain 3D objects from OmniObject3D, Objaverse, and Replica, which include both house and non-house categories.
@@
Category Dataset NBV Policy AUC $\uparrow$ FCR $\uparrow$ Acc. $\downarrow$
@@
Implementation Details. We conduct all experiments in Isaac Gym simulation engine with one NVIDIA Tesla V100 GPU. Our NBV policy is optimized through over 32 million iterations and uses approximately 24 hours of training time on an NVIDIA V100 GPU. All networks are randomly initialized and trained end-to-end. Our implementation refers to the codebase of Legged Gym [33] and the PPO implementation in Stable Baseline3 [29], which is implemented by PyTorch [53]. The ground-truth point clouds on objects’ surfaces are sampled by the Poisson Disk sampling method [47] using Open3D API [51]. The Chamfer Distance between the scanned point cloud and the ground-truth point cloud is calculated using PyTorch3D API [31]. Please refer to the Appendix for further details and experiments.
@@ 4.2 Performance Comparison
Non-house Generalization In Table 2, we introduce an experiment to evaluate the generalizability of our GenNBV trained on Houses3K to non-house categories. The targets include $74$ animals, $43$ trucks and $33$ dinosaurs from OmniObject3D, and a challenging indoor scene (Room 0) from Replica. The number of views is set to 20.
@@ 4.3 Ablation Study
We implement the ablation study about multi-source state embedding and depth-based representation. Please refer to the Appendix for further experiments.
@@
4.4 Qualitative Results
5 Conclusion
This study presents an end-to-end approach for active 3D scene reconstruction, reducing the need for manual intervention. Specifically, the learning-based policy explores how to reconstruct diverse objects in the training stage and thus can generalize to reconstruct unseen objects in a fully autonomous manner. Our controller maneuvers in free space and selects the next best view based on a hybrid scene representation which conveys scene coverage status and thus reconstruction progress. We show the effectiveness of our approach by generalizing it on multiple datasets. The quantitative and qualitative generalization results on holdout Houses3K test set and cross-domain OmniObject3D, including house category and non-house categories, show that our method outperforms other baselines in terms of reconstruction completeness, efficiency and accuracy. Furthermore, the experiment conducted on Objaverse and Replica shows that the policy trained in house-only settings can even generalize to complicated outdoor and indoor scenes.
Acknowledgements
This work is supported by Shanghai Artificial Intelligence Laboratory and CUHK Direct Grants (RCFUS) 4055189. We’re very grateful to Tao Huang, Zeqi Xiao, Junfeng Long, Mulin Yu, and all reviewers for their insightful advice.
Supplementary Material
Appendix A Implementation Details
A.1 Occupancy Grid Mapping Algorithm
. Occupancy Threshold Mean AUC $\uparrow$ Final Coverage Ratio $\uparrow$ Accuracy $\downarrow$ 0.5 (Ours) 87.39% 95.63% 0.37 1.0 84.41% 93.43% 0.41 1.5 84.84% 95.20% 0.39 2.5 54.77% 58.70% 0.85 Value of Constant $C^{\prime}$ Mean AUC $\uparrow$ Final Coverage Ratio $\uparrow$ Accuracy $\downarrow$ 5 66.96% 71.92% 0.49 10 80.75% 92.16% 0.43 20 (Ours) 87.39% 95.63% 0.37 40 86.97% 94.12% 0.40
A.2 Network and Training
Training. In Isaac Gym [22], we set 256 parallel environments for policy training, where each environment corresponds with one building-level object. Once the coverage ratio reaches a threshold (99% in practice), a collision happens or the episode length reaches the maximum threshold (100 steps), the environment is reset and the building to be reconstructed is replaced. Our NBV policy is optimized through over 32 million iterations and uses approximately 24 hours of training time on an NVIDIA V100 GPU. All networks are randomly initialized and trained end-to-end.
$s_{t}=\mathrm{Linear}(s^{G}_{t};s^{S}_{t};s^{A}_{t}),$ (7)
where all embeddings are 256 vectors.
Our policy network PPO, implemented by Stable Baseline3 [29], is a 3-layer multi-layer perceptron (MLP). The output of our policy is used to parameterize a distribution over our 5-dimension action space. In this way, the action can be drawn from the stochastic policy $a\sim\pi(\cdot|o_{t})$.
A.3 Baseline Policies
Appendix B Data Preprocessing
Appendix C Additional Experiments
@@
The effectiveness of 2D semantic representations. Below, we add the experiments on Houses3K test set to illustrate the effectiveness of the proposed semantic representation. Note that Scan-RL’s representation uses 6 frames, while our semantic representation only needs 2 frames.
Num. of RGB NBV Policy AUC $\uparrow$ FCR $\uparrow$ Acc. $\downarrow$